I have to create a system of equations from the given word problem and I can't figure out how to do it... the word problem is... As a participant in your schools community service project, you volunteer a total of 40 hours over the course of the school year. Your volunteer hours include serving at a soup kitchen, picking up trash at several local parks, and collecting toys for needy children. You spend 4 times as many hours collecting toys as picking up trash, and 2 hours less serving at the soup kitchen than picking up trash. Write a system of equations to represent the given information. Define your variables as s,p, and c for the hours of serving at a soup kitchen, picking up trash, and collecting toys. Can you help me please?
Answer by Edwin McCravy(20086)

I have to create a system of equations from 
the given word problem and I can't figure out 
how to do it... the word problem is... As a 
participant in your schools community service 
project, you volunteer a total of 40 hours 
over the course of the school year. Your 
volunteer hours include serving at a soup 
kitchen, picking up trash at several local 
parks, and collecting toys for needy children. 
You spend 4 times as many hours collecting toys 
as picking up trash, and 2 hours less serving at 
the soup kitchen than picking up trash. Write a 
system of equations to represent the given 
information. Define your variables as s,p, and c 
for the hours of serving at a soup kitchen, 
picking up trash, and collecting toys. Can you 
help me please?

>>...a total of 40 hours...<<

Translation:  s + p + c = 40

>>...4 times as many hours collecting 
      toys as picking up trash...<<

Translation:          c = 4p

>>...2 hours less serving at the soup 
      kitchen than picking up trash...<<

Translation:          s = p - 2


So the system of equations is:

s + p + c = 40
        c = 4p
        s = p - 2
 
Can you solve that system?  If not post again 
asking how.

Answer:  s = 5 hr, p = 7 hr, c = 28 hr

Edwin